In rectangle ABCD, diagonals AC and BD intersect at Point E. If AC=36 and BD=2x+30, find x.
Answer by math-vortex(648)

Hi, there--

THE PROBLEM:
In rectangle ABCD, diagonals AC and BD intersect at Point E. If AC=36 and BD=2x+30, find x.

A SOLUTION:
To solve this problem, we use the mathematical fact the the diagonals of a rectangle are 
equal in length. Therefore the length of AC is equal to the length of BD.

We are given that the length of AC is 36, and the length of BD is 2x+30. We set these equal 
to each other and solve for x.

36 = 2x + 30
2x = 36 - 30
2x = 6
x = 3

If x is 3 then the length of BD is 
2x + 30 = 2(3) + 30 = 36

Both diagonals at 36 units in measure. CHECK!

Hope this helps!

Mrs. Figgy
math.in.the.vortex@gmail.com